OpenStudy (anonymous):

10x-5y=-20 7x-5y=-26

OpenStudy (vickyisthesmartone123):

I know the answer hold on...

OpenStudy (vickyisthesmartone123):

(2,8) That is the answer!:)

OpenStudy (anonymous):

Try and use substitution. Here is an example: The substitution method is a method that is based off a rule of algebra. Basically, it's similar to using a formula. You substitute values into the formula. The same thing applies here. Let's say we have the following system: y = 2x + 3 5x + 2y = 24 Before we turn over to elimination, let's try using our new friend substitution. It clearly says that y is equal to 2x + 3. This means that we can substitute 2x + 3 in for y. 5x + 2y = 24 5x + 2(2x + 3) = 24 5x + 4x + 6 = 24 Now that we have eliminated a variable, we can solve the system. 9x + 6 = 24 9x = 18 x = 2 Now let's plug this back into one of the equations. I prefer to plug it into the equation that has a variable isolated. y = 2x + 3 y = 2(2) + 3 y = 4 + 3 y = 7 Therefore, the solution set is (x, y) = (2, 7).

OpenStudy (vickyisthesmartone123):

No it's not it's (2,8)

OpenStudy (anonymous):

See if you understand substitution, and try and use it on your own equations :)

OpenStudy (vickyisthesmartone123):

You copied and pasted that im not dumb

OpenStudy (vickyisthesmartone123):

From another person

OpenStudy (vickyisthesmartone123):

You can't type that fast

OpenStudy (tygrr321):

OOOHHHHHH the suspense

OpenStudy (anonymous):

Super fast typer!

OpenStudy (tygrr321):

jgdskgfuilfehhywsuyhjhdfilugiaupiojrgijeirsjil;jhilr;ji;ujgijhrjgs;kjlisuilpj;/auilrauululerhgilrhiulgujikdfsjkigujerijgikdjikdfjifjkjire irjgivjliksjro;jger;jr erhieworggjerjilg;er goegleroergr rekgerj;gkfldjkod;ikrkhyjeopkiyhp'ete rgeierjgejgkl;eg er eer;oergjrgt;jerje rejkhrngklejh;kgj;erkj;gljk;kj'olrgk'erjrk'hg'olerh'kehr

OpenStudy (anonymous):

But it doesnt matter, it's all about making him understand how to use the substitution method.

OpenStudy (vickyisthesmartone123):

It taes so long just to get to the point!

OpenStudy (vickyisthesmartone123):

*takes

OpenStudy (anonymous):

But he needs to learn it first, you can't learn something new in 2 seconds.

OpenStudy (tygrr321):

only if its 15 seconds than yeah

OpenStudy (vickyisthesmartone123):

BYE COPYING AND PASTING? SHEESH

OpenStudy (vickyisthesmartone123):

YOUR THE ONE THAT'S NOT GOING TO LEARN

OpenStudy (anonymous):

? It doesnt matter where I found the text, if it helps him learn.

OpenStudy (anonymous):

I dont need to learn anything from this post, I am not the one who asked it.

OpenStudy (vickyisthesmartone123):

bye